For the curves y=x^2+4x+8 and y=x^2+8x+4,the equation of common tangent is
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
for first parabola
 
(y- 4) = (x+2)^2
 
now
equation of tangent
 
( y- 4) = m (x+4) + 1/m
 
now putvthis equation in 2nd hyperbola and make discriminant zero then you will get value of m.
 
then you will clearly get the equation of common tangent